### v2

# Merge Kernel for Bayesian Optimization on Permutation Space

Abstract
Bayesian Optimization (BO) algorithm is a standard tool for black-box optimization problems. The current state-of-the-art BO approach for permutation spaces relies on the Mallows kernel-an $\Omega(n^{2})$ representation that explicitly enumerates every pairwise comparison. Inspired by the close relationship between the Mallows kernel and pairwise comparison, we propose a novel framework for generating kernel functions on permutation space based on sorting algorithms. Within this framework, the Mallows kernel can be viewed as a special instance derived from enumeration sort. Further, we introduce the Merge Kernel constructed from merge sort, which replaces the quadratic complexity with $\Theta(n\log n)$ to achieve the lowest possible complexity. The resulting feature vector is significantly shorter, can be computed in linearithmic time, yet still efficiently captures meaningful permutation distances. To boost robustness and right-invariance without sacrificing compactness, we further incorporate three lightweight, task-agnostic descriptors: (1) a shift histogram, which aggregates absolute element displacements and supplies a global misplacement signal; (2) a split-pair line, which encodes selected long-range comparisons by aligning elements across the two halves of the whole permutation; and (3) sliding-window motifs, which summarize local order patterns that influence near-neighbor objectives. Our empirical evaluation demonstrates that the proposed kernel consistently outperforms the state-of-the-art Mallows kernel across various permutation optimization benchmarks. Results confirm that the Merge Kernel provides a more compact yet more effective solution for Bayesian optimization in permutation space.
Introduction
As one of the most widely adopted approaches to black-box optimization, Bayesian optimization (BO) (Shahriari et al. 2015) has found broad application in machine-learning hyper-parameter tuning (Wu et al. 2019), financial portfolio optimization (Gonzalvez et al. 2019), chemical‐material discovery (Luo et al. 2025), and catalyst formulation design (Xie et al. 2023). BO employs probabilistic surrogate models—most commonly Gaussian processes (GPs)—to approximate the unknown objective and uses an acquisition function to balance exploration and exploitation, thereby approaching the global optimum with the fewest possible evaluations.
Most research on BO to date has concentrated on continuous (Greenhill et al. 2020) and categorical design spaces (Garrido-Merchán and Hernández-Lobato 2020; Nguyen et al. 2020; Bartoli et al. 2025), whereas applications to permutation spaces remain comparatively underexplored. This gap is striking given the ubiquity of permutation optimization in both theory and practice: canonical examples include the travelling-salesperson problem (TSP), the sequencing of operations in automated experimental pipelines (Guidi, Seeberger, and Gilmore 2020), and the sequential order-of-addition experiments (Lin and Rios 2025). In addition, permutation optimization frequently arises in diverse AI applications, including scheduling tasks in robotic planning (Alatartsev, Stellmacher, and Ortmeier 2015), optimizing experimental sequences (Blau et al. 2022), and other sequential decision-making problems (Sun and Giles 2001; Wen et al. 2023). Therefore, extending Bayesian optimization to permutation spaces holds significant theoretical and practical value for artificial intelligence as well as broader academic and industrial applications.
Successfully deploying BO in permutation spaces hinges on equipping the GP surrogate with a kernel that faithfully quantifies the similarity between two permutations. Existing approaches have evolved along two principal lines: (1) General-purpose discrete BO frameworks—most notably COMBO (Oh et al. 2019). COMBO combines a graph-Laplacian kernel with local search to accommodate heterogeneous discrete variables and thus enjoys broad applicability. However, it relies on manually crafted adjacency graphs or hash encodings, which struggle to capture the comparison sequences and cyclic shifts unique to permutations; consequently, its efficiency degrades on tasks that involve frequent high-order swaps or mixed local/global reorderings. (2) Permutation-specific kernels—chief among them the Mallows kernel in BOPS(Deshwal et al. 2022). This kernel is built on the Kendall–$\tau$ distance, representing a permutation via the inversion counts generated by enumeration sort, yielding a feature dimension of $O(n^{2})$. The dimension therefore scales quadratically with $n$, and because $2^{n^{2}}\!\gg n!$, the vast majority of features fail to correspond to any valid permutation, resulting in both statistical and computational redundancy.
To address the limitations of existing approaches, we propose a sorting-algorithm–driven kernel-design framework for permutations and instantiate it with the Merge kernel, which reduces the feature dimension to $O(n\log n)$—the information-theoretic lower bound for encoding a permutation. The central insight is that any comparison-based sorting algorithm is defined by a fixed sequence of element comparisons; recording the binary outcome of each comparison yields a feature vector for the permutation. Choosing an algorithm with a deterministic comparison tree—such as merge sort or bitonic sort—thus produces a representation that is both fixed in length and highly compact.
To further enhance robustness and expressiveness, we augment the Merge kernel with three lightweight descriptors:
  • Shift Histogram – captures invariance under cyclic shifts;
  • Split-Pair Line – emphasises the relative order of elements across partition boundaries;
  • Sliding-Window Motif – extracts local sorting patterns, increasing sensitivity to local perturbations.
This design attains the shortest possible encoding while simultaneously eliminating the high-dimensional redundancy of the Mallows kernel and the structural blind spots of COMBO, thereby providing an efficient and accurate similarity measure for Bayesian optimization over permutation spaces.
Contributions. Our work makes three principal contributions:
  • General framework. We propose a unified design framework that constructs permutation-space kernels by treating any comparison-based sorting algorithm as a feature generator. Within this view, the classic Mallows kernel is recovered as the special case obtained when the framework is instantiated with enumeration sort.
  • Merge kernel. Applying the framework to merge sort yields Merge Kernel, whose $O(n\log n)$ construction matches the information-theoretic lower bound on comparison complexity. We further enrich it with three lightweight descriptors to boost predictive power at minimal cost.
  • Comprehensive evaluation. We assess the effectiveness of our kernels on diverse synthetic and real-world benchmarks, consistently outperforming the state-of-the-art Mallows kernel in both accuracy and efficiency.
Our results demonstrate that the Merge kernel provides a practical and efficient tool for permutation optimization, significantly enhancing BO’s applicability to diverse AI scenarios.
Background and Related Works
Permutation Optimization
Here we describe the problem formulation of permutation optimization with a fixed length $n\in\mathcal{N}$. Let $[n]=\{1,2,...,n\}$, a permutation is a function $\pi:[n]\longrightarrow[n]$ such that $\pi$ is bijective. The set of all permutations of $[n]$ is the symmetric group
$\mathcal{S}_{n}\;=\;\bigl{\{}\pi\mid\pi:[n]\longrightarrow[n]\text{ is bijective}\bigr{\}}.$
We are given a costly-to-evaluate, possibly noisy black-box function
$f:\;\mathcal{S}_{n}\longrightarrow\mathbb{R},$
which assigns a real-valued quality (e.g., cost, loss, reward) to every permutation $\pi$. Hence, the optimization problem can be formulated as
$\pi^{*}\;=\;\text{arg min}_{\pi\;\in\;\mathcal{F}}f(\pi)$
where $\mathcal{F}\subseteq\mathcal{S}_{n}$ is the feasible set. In this study we only consider the unconstrained case, therefore we have $\mathcal{F}\;=\;\mathcal{S}_{n}$; in practice $\mathcal{F}$ may exclude permutations violating domain rules.
Bayesian Optimization
Bayesian optimization (BO) (Shahriari et al. 2015) is an optimization algorithm for black-box objective functions that no closed-form expression or gradient information is available and whose evaluation is often an expensive physical or computational experiment. The algorithm first fits the observed data with a surrogate model, most commonly a Gaussian process (GP) (Williams and Rasmussen 2006), and then employs an acquisition function to select the next query point, balancing exploration of uncertain regions against exploitation of promising areas identified by the surrogate.
The kernel $K(\mathbf{x},\mathbf{x}^{\prime})$ is the central design lever in a Gaussian-process surrogate: it defines the similarity metric between inputs, thereby specifying the prior smoothness assumptions and, through GP inference, the posterior mean and uncertainty. Extending BO to any new search domain is therefore tantamount to endowing that domain with an appropriate kernel function. Whereas Euclidean spaces typically rely on Gaussian (RBF) kernels, discrete structures—and permutations in particular—require bespoke constructions that faithfully encode ordering relationships. In the next subsection we revisit the Mallows kernel on the symmetric group $\mathcal{S}_{n}$, outline its computational and statistical shortcomings, and use these observations to motivate our own kernel design.
Mallows Kernel for Permutation Space
BOPS-H (Deshwal et al. 2022) is the current state-of-the-art BO algorithm for permutation optimization, which propose to employ Mallows kernel (Jiao and Vert 2015) on the symmetric group $\mathcal{S}_{n}$ in a similar manner to the RBF kernel on the Euclidean space. The Mallows kernel $K_{mal}(\pi,\pi^{\prime})$ for the permutation pair ($\pi$, $\pi^{\prime}$) is defined as the exponential negative of the number of discordant pairs $n_{d}(\pi,\pi^{\prime})$ between $\pi$ and $\pi^{\prime}$:
$K_{Mal}(\pi,\pi^{\prime})=\exp(-ld(\pi,\pi^{\prime}))$ (1)
where $l\geq 0$ is the length-scale parameter of the Mallows kernel, and $d(\pi,\pi^{\prime})$ is the Kendall-$\tau$ distance (Kendall 1938) which counts the number of pairs of elements ordered oppositely by $\pi$ and $\pi^{\prime}$:
$d\left(\pi,\pi^{\prime}\right)=\sum_{i<j}$ $\left[1_{\pi(i)>\pi(j)}1_{\pi^{\prime}(i)<\pi^{\prime}(j)}\right.$ (2)
$\left.\quad+1_{\pi(i)<\pi(j)}1_{\pi^{\prime}(i)>\pi^{\prime}(j)}\right]$
Intuitively, the Kendall-$\tau$ distance counts the differences of all pair-wise comparisons between $\pi$ and $\pi^{\prime}$. For example, let $\pi\;=\;(1,2,3,4)$ and $\pi^{\prime}\;=\;(2,1,4,3)$. Two pairs are discordant among the six unordered pairs: $(1,2),(3,4)$, hence $d(\pi,\pi^{\prime})=2$ and $K_{Mal}(\pi,\pi^{\prime})=\exp(-2l)$.
Algorithm
Generating Kernel based on Sorting Algorithm
In the previous section, we have shown that the core of the Mallows kernel is the pairwise comparison of all elements. Equivalently, it maps a permutation $\pi$ to a feature vector
$\Phi_{Mal}(\pi)\;\in\;\{-1,1\}^{\binom{n}{2}}$
where each coordinate corresponds to the comparison of a pair of elements: -1 if they are in ascending order, and +1 otherwise. We then have
$\begin{split}K_{Mal}(\pi,\sigma)&=\exp\!\Bigl{(}-\frac{\|\Phi_{Mal}(\pi)-\Phi_{Mal}(\sigma)\|^{2}}{2\ell^{2}}\Bigr{)}\\ &=\;k_{\mathrm{RBF}}\bigl{(}\Phi_{Mal}(\pi),\,\Phi_{Mal}(\sigma)\bigr{)}.\end{split}$ (3)
Since the Gaussian RBF kernel $k_{\mathrm{RBF}}$ is strictly positive definite on $\mathbb{R}^{d}$ and thus satisfies Mercer’s condition (Mercer 1909), and because positive definiteness is preserved under composition with any deterministic mapping $\Phi$, it follows that $K(\pi,\sigma)$ constructed from $\Phi$ also satisfies Mercer’s condition and is therefore a valid kernel function.
Thus, other pairwise comparison methods can also be used to construct analogous feature vectors, which—when combined with an RBF kernel—yield valid kernel functions. Naturally, we can extend the idea of pairwise comparison to sorting algorithms: the essence of a sorting algorithm is to compare elements in a sequence and swap them when necessary. Consequently, each sorting algorithm embodies a unique pairwise comparison strategy, suggesting that we can build permutation‐space kernels based on sorting procedures. As a sorting algorithm traverses all elements, it records whether each comparison leads to a swap, thereby fully reconstructing the original permutation; hence, the resulting feature vector retains all information without any loss. Viewed in this light, the Mallows kernel’s exhaustive enumeration of every element pair can be interpreted as an enumeration-sort–inspired featurization, where enumeration sort ranks each item by comparing it with every other element and then places it directly in its final position.
However, not every sorting algorithm can induce a valid feature mapping suitable for kernel construction. This is because the mapping from permutations to feature space must yield feature vectors of fixed length; otherwise, feature vectors of differing lengths would not be compatible with the RBF kernel. Hence, only sorting algorithms that have a fixed comparison path and a constant number of comparisons across all inputs can generate valid feature mappings. In fact, most sorting algorithms commonly known to have complexity of order $O(n\log n)$—such as quicksort, heapsort, and standard merge sort—do not satisfy the requirement of constant comparison counts. Nevertheless, merge sort is an exception: although typically merge sort ceases comparisons once all elements from one subsequence have been merged, redundant comparisons can be artificially introduced during the merge procedure. This ensures that both the comparison path and the number of comparisons remain identical across different permutations, thus establishing a fixed comparison path. Consequently, we select merge sort as the feature mapping $\Phi_{Mer}(\pi)$ for constructing permutation-space kernel function $K_{Mer}$, following Equation 3:
$\begin{split}K_{Mer}(\pi,\sigma)&=\exp\!\Bigl{(}-\frac{\|\Phi_{Mer}(\pi)-\Phi_{Mer}(\sigma)\|^{2}}{2\ell^{2}}\Bigr{)}\\ &=\;k_{\mathrm{RBF}}\bigl{(}\Phi_{Mer}(\pi),\,\Phi_{Mer}(\sigma)\bigr{)}.\end{split}$ (4)
The element-pair comparison mapping $\Phi_{Mer}(\pi)$ is shown in Algorithm 1.
Algorithm 1 Merge feature mapping $\Phi_{Mer}(\pi)$
Input: Permutation $\pi$ of length $n$
Output: Feature vector $\Phi_{Mer}(\pi)$
  if length($\pi$) == 1 then
     return []
  end if
  if length($\pi$) == 2 then
     return [1] if $\pi[0]>\pi$[1]
     return [-1] if $\pi[0]<\pi$[1]
  end if
  Let $mid=\lfloor\frac{n}{2}\rfloor$
  Let $V_{Left}=\Phi_{Mer}(\pi[:mid])$
  Let $V_{Right}=\Phi_{Mer}(\pi[mid:])$
  Let $\hat{\pi}_{l}=\text{sorted}(\pi[:mid])$
  Let $\hat{\pi}_{r}=\text{sorted}(\pi[mid:])$
  Let $V_{Merge}=[]$
  Let $i=j=0$
  while $i<\text{length}(\hat{\pi}_{l})$ and $j<\text{length}(\hat{\pi}_{r})$ do
     if $\hat{\pi}_{l}[i]>\hat{\pi}_{r}[j]$ then
        $V_{Merge}$.append(1)
        $j+=1$
     end if
     if $\hat{\pi}_{l}[i]<\hat{\pi}_{r}[j]$ then
        $V_{Merge}$.append(-1)
        $i+=1$
     end if
  end while
  return $V_{Left}+V_{Right}+V_{Merge}$
We have established that merge sort, with a specially designed fixed-comparison procedure, is uniquely capable of constructing valid kernel functions among common sorting algorithms with a complexity of $\Omega(n\log n)$. We now demonstrate that the feature vector derived from merge sort achieves the theoretical lower bound on vector length for lossless permutation encoding. First, note that the lower bound on time complexity for any comparison-based sorting algorithm is $\Omega(n\log n)$; as this complexity directly corresponds to the number of element comparisons during sorting, it similarly sets a lower bound on the length of the feature vector. On the other hand, consider the permutation space consisting of all $n!$ possible permutations of length $n$. From an information-theoretic viewpoint, encoding all $n!$ permutations without loss using a binary feature vector composed solely of ${-1,1}$ requires a minimum vector length of $\log_{2}(n!)$. Applying Stirling’s approximation (Donald et al. 1999), we have:
$\log_{2}(n!)=n\log_{2}n-nlog_{2}e+O(\log n)=\Omega(n\log n)$
Consequently, the feature vector length of the merge-sort-based kernel (Merge Kernel) reaches this theoretical lower bound for lossless permutation encoding.
Adding Invariance to Merge Kernel
It is worth noting that, when constructing the Merge kernel via merge sort, we have not required the feature mapping $\Phi$ to possess any group invariance property, such as right-invariance. Traditionally, a permutation-distance measure should be invariant under right multiplication, meaning that applying an identical right-multiplication operation to two permutations should not alter the distance between them. However, only sorting algorithms with complexity $O(n^{2})$ can yield fully right-invariant kernel functions, since such invariance necessitates exhaustive pairwise comparisons among all $\frac{n(n-1)}{2}$ pairs of elements—an impossibility for more efficient sorting algorithms like merge sort. Consequently, although the Merge kernel achieves better computational efficiency through a more compact encoding, it sacrifices a certain degree of performance due to the loss of right-invariance.
To address this issue, we can naturally consider trading off some compactness by slightly increasing the encoding length—while preserving the overall $\Omega(n\log n)$ complexity—to enhance right-invariance and thus compensate for the associated performance loss. We introduce three additional encoding strategies, whose resulting feature vectors are concatenated to the original Merge feature mapping to further improve the kernel’s performance.
Algorithm 2 Sliding-window Motif Feature Mapping $\Phi_{Sli}(\pi)$
Input: permutation $\pi$ of length $n$; window length $w$
Output: sliding-window motif vector $\Phi_{Sli}(\pi)$
[…]
  build dictionary $\text{index}[p]\leftarrow$ rank of pattern $p$ in patterns
  $H\leftarrow\mathbf{0}_{\,w!}$initialise histogram}
  for $i\leftarrow 0$ to $n-1$ do
     $\text{win}\leftarrow\bigl{(}\pi_{(i+j)\bmod n}\bigr{)}_{j=0}^{w-1}$circular slice}
     $p\leftarrow\textsc{Relative Order}(\text{win})$
     $idx\leftarrow\text{index}[p]$
     $H[idx]\leftarrow H[idx]+1$
  end for
  return $H\;\;(=\Phi_{Sli}(\pi))$
Algorithm 3 Relative Order$(v)$
Input: sequence $v=(v_{0},\dots,v_{w-1})$
Output: permutation $p$ giving the relative order of $v$
[…]
  $p\leftarrow$ array of length $w$
[…]
     $p[i]\leftarrow r$
  end for
  return tuple$(p)$
Algorithm 4 Shift-histogram feature mapping $\Phi_{Shi}(\pi)$
Input: Permutation $\pi$ of length $n$, maximum shift $s_{\max}$
Output: Shift-histogram feature vector $\Phi_{Sli}(\pi)$
  $n\leftarrow\text{length}(\pi)$
  for $i\leftarrow 0$ to $n-1$ do
     $\text{pos}[\pi[i]]\leftarrow i$
  end for
  for $i\leftarrow 0$ to $n-1$ do
     $\Delta_{i}\leftarrow\text{pos}[i]-i$
     if $\Delta_{i}>s_{\max}$ then
        $\Delta_{i}\leftarrow s_{\max}$
     else if $\Delta_{i}<-s_{\max}$ then
        $\Delta_{i}\leftarrow-s_{\max}$
     end if
  end for
[…]
  for $i\leftarrow 0$ to $n-1$ do
     $H[\Delta_{i}+s_{\max}]\leftarrow H[\Delta_{i}+s_{\max}]+1$
  end for
  return $H$
Middle-split cross-column pair comparison
Although merge sort efficiently encodes comparisons within local subsets, it does not explicitly cover cross-partition pairs at each recursive split, specifically the first split—the comparison vector $V_{Merge}$ will be computed in the last recursion after left half and right half sorted. To remedy this and capture long-range interactions, we propose to add the strict pairwise comparisons $\Phi_{Mid}(\pi)$ between elements sharing the same offset in their respective halves after first split. For example, $\Phi_{Mid}(\pi)$ will have 3 digits when $\pi$ has length of 6: comparison between $(\pi_{0},\pi_{3}),(\pi_{1},\pi_{4}),(\pi_{2},\pi_{5})$. To keep the representation compact, we retain only the cross-partition comparisons produced by the outermost split. At this level the two halves are merged only in the final recursion step, so their relative order is maximally perturbed and the partition gap is the largest, encapsulating the richest set of inter-partition comparisons while requiring only $\Omega(n)$ encoding length. Because the resulting procedure is straightforward, we omit its pseudocode here.
Sliding-window motif
To capture local ordering patterns that heavily influence near-neighbor-dependent objectives and add cyclic-rotation invariance, we introduce sliding-window motifs. For a small number of window length $k$, we enumerate all possible local ordering patterns (motifs) and count their occurrences within the permutation. With three-element windows, there are exactly $k!$ distinct relative-order patterns. We store their frequencies as a compact local-order texture descriptor, further enriching the kernel with information on short-range permutation structures. Because the sliding window wraps around and fills from the beginning of the permutation once it reaches the end, this scheme explicitly endows the features with cyclic-rotation invariance. The feature mapping algorithm $\Phi_{Sli}(\pi)$ is shown in Algorithm 2 and 3.
Shift-histogram
To incorporate another global misplacement measure that remains invariant under cyclic right-multiplication, we introduce the shift histogram $\Phi_{Shi}$. For each permutation $\pi$, we first calculate the absolute displacement of each element relative to its position in the identity permutation, defined formally as:
$d_{i}=\pi^{-1}(i)-i,\forall i\in\{0,1,...,n-1\}$
The displacements are then binned into a fixed-size histogram, typically truncated at some maximum absolute shift (e.g.,±$\pm 5$), resulting in a compact, stable global descriptor of permutation misplacement. The feature mapping algorithm $\Phi_{Shi}(\pi)$ is shown in Algorithm 4. Eventually, the final Merge kernel uses a concatenated feature mapping strategy contains all elements above:
$\Phi_{Merge}(\pi)=concat(\Phi_{Mer}(\pi),\Phi_{Mid}(\pi),\Phi_{Sli}(\pi),\Phi_{Shi}(\pi))$
Intuitively, these lightweight descriptors enhance the kernel’s ability to capture global and local permutation structures, improving both optimization stability and robustness.
Experiments
Figure 1: Results comparing Mallows kernel (BOPS-H) and Merge kernel (MergeBO) on the current regret value (difference between best-so-far and optimal value) vs. number of iteration.
Simple final regret Best so far AUC
Merge Mallows Merge Mallows
$\text{TSP}_{n=15}$ 0.012$\pm$0.038 0.012$\pm$0.038 400.9$\pm$115.0 428.2$\pm$121.9
$\text{QAP}_{n=15}$ 6578.0$\pm$4579.2 8090.4$\pm$4101.9 2493028.7$\pm$690320.7 2750160.0$\pm$744154.9
$\text{FloorPlanning}_{n=30}$ 23.3$\pm$11.5 30.1$\pm$12.8 8140.8$\pm$2386.8 8665.7$\pm$2638.8
$\text{CellReplacement}_{n=30}$ 6.1$\pm$2.1 6.1$\pm$3.0 5371.0$\pm$792.7 5350.5$\pm$910.1
Table 1: Performance comparison between MergeBO and BOPS-H (Mallows kernel), better performance shown in bold font
Simple final regret All tricks No pairwise No pattern No shift
$\text{TSP}_{n=15}$ 0.012$\pm$0.038 0.025$\pm$0.051 0.038$\pm$0.059 0.019$\pm$0.046
$\text{QAP}_{n=15}$ 6578.0$\pm$4579.2 6330.0 $\pm$3132.2 6428.2$\pm$3777.5 8280.6$\pm$4536.0
$\text{FloorPlanning}_{n=30}$ 23.3$\pm$11.5 25.4$\pm$9.4 26.9$\pm$11.5 26.2$\pm$10.5
$\text{CellReplacement}_{n=30}$ 6.1$\pm$2.1 5.1$\pm$2.1 6.1$\pm$2.4 6.3$\pm$3.7
Best so far AUC All tricks No pairwise No pattern No shift
$\text{TSP}_{n=15}$ 400.9$\pm$115.0 418.7$\pm$125.3 471.6$\pm$167.5 410.8$\pm$107.3
$\text{QAP}_{n=15}$ 2493028.7$\pm$690320.7 2588481.3$\pm$638766.7 2580381.5$\pm$838626.7 2975522.9$\pm$743481.9
$\text{FloorPlanning}_{n=30}$ 8140.8$\pm$2386.8 8311.2$\pm$2202.0 7940.6$\pm$1908.8 8379.9$\pm$2237.7
$\text{CellReplacement}_{n=30}$ 5371.0$\pm$792.7 5553.1$\pm$1096.4 5390.3$\pm$952.2 5590.6$\pm$1040.5
Table 2: Ablation experiment results with each trick dropped, better performance shown in bold font
In this section, we demonstrate the benchmarks, the experimental settings, the ablation studies as well as the results and discussion.
Benchmarks and Experiment Settings
Because our method, MergeBO, is a direct generalisation of the state-of-the-art Mallows kernel in BOPS-H algorithm(Deshwal et al. 2022), we restrict the empirical evaluation to that baseline and adopt the identical benchmark suite and experimental protocol introduced in (Deshwal et al. 2022). This isolates the effect of replacing the Mallows kernel with our merge-based variant and eliminates confounding factors due to data splits or hyper-parameter choices.
The BOPS-H algorithm follows and modifies the local-search strategy used in COMBO, examining only the set of neighbouring permutations of the current incumbent to restrict the combinatorial search space, we therefore adopt the same procedure in our experiments. To keep the extra features introduced by the tricks compact and prevent them from overshadowing the Merge kernel, we set the sliding‑window length to 4 and capped the shift‑histogram at a maximum shift of 5 for all four tasks. GPyTorch (Gardner et al. 2018) and BoTorch (Balandat et al. 2020) libraries are used to implement both alogithms. We evaluate our method on the same two synthetic benchmarks and the same two real-world applications in (Deshwal et al. 2022). Detailed information for all benchmarks are listed below:
(1)Quadratic Assignment (QAP). This is a classic facility-location problem: assign $n$ facilities to $n$ locations so that flow costs and distances align optimally. We use the 15-city instances from QAPLIB (Burkard, Karisch, and Rendl 1997), each with cost matrix $A$ and distance matrix $B$, and minimise $\operatorname{Tr}(APBP^{\mathsf{T}})$ over permutation matrices $P$.
(2)Travelling Salesman (TSP). The TSP seeks the shortest Hamiltonian cycle through a set of cities and is a standard benchmark for route-planning. Our instances are ten-city PCB drill tours from TSPLIB (Reinelt 1995); the score is the total travel time to visit all holes exactly once and return.
(3)Floor Planning (FP). Floor planning is an NP-hard VLSI layout task that packs rectangular modules on a chip while minimising area and manufacturing cost. We evaluate two ten-block variants (FP-1 and FP-2); each permutation defines a block placement whose cost we minimise.
(4)Cell Placement (CP). Cell placement arranges logic cells on a row to reduce wire-length and hence circuit delay.We consider ten equal-height cells with a fixed net-list; the objective is the total Manhattan wire-length induced by a permutation of cell positions.
Because the publicly available implementation of the Mallows kernel (https://github.com/aryandeshwal/BOPS) does not provide the interface required to run the Rodinia’s heterogeneous many-core benchmark (Che et al. 2009), we did not perform experiments on this benchmark.
Expected Improvement acquisition function is used for all the experiments, and 10 restarts are used for local search based EI optimization for BOPS-H and MergeBO. In this study, we employ two evaluation metrics: the final simple regret and the area under the best‑so‑far regret curve (AUC). In optimization, regret is defined as the difference between the best objective value observed to date and the global optimum:
$r_{t}=f_{t}^{\text{best}}-f^{*}$
Following this concept, the final simple regret is the regret value obtained in the last iteration and reflects the algorithm’s ultimate optimization capability when computational cost is disregarded. By contrast, the regret AUC—the sum (area under the curve) of the simple regret across all iterations—quantifies the convergence speed of the entire optimization process:
$\text{AUC}_{T}=\Sigma r_{t}$
All experiments were executed for 200 iterations and repeated 20 times using an identical sequence of random seeds. Algorithm performance was evaluated by the mean and variance of the final simple regret and regret AUC on all 20 trials; in both cases, smaller values indicate better performance.
Results and Discussion
In this subsection, we present and discuss our experimental results along different dimensions against current SOTA algorithm BOPS-H.
As shown in figure 1, the Merge kernel consistently outperforms the Mallows kernel across all four permutation optimization benchmarks. The regret curves demonstrate that MergeBO achieves faster convergence and lower mean regret compared to BOPS-H throughout the entire optimization process. Notably, the advantage is particularly pronounced on more challenging tasks such as QAP and Floor Planning, where the Merge kernel exhibits both reduced variance and a more rapid decrease in regret. The TSP problem is a relatively common and tractable combinatorial optimization problem, so both algorithms are able to approach the global optimum in the end. Nevertheless, it is evident that the Merge kernel achieves a faster convergence rate compared to the Mallows kernel. These results highlight the improved sample efficiency and robustness of the Merge kernel in complex permutation spaces.
Table 1 provides a quantitative summary of performance, reporting both the final simple regret and the best-so-far area under the curve (AUC) for each benchmark. Merge kernel achieves either the lowest or statistically equivalent regret on all tasks, and outperforms Mallows kernel in three out of four benchmarks in terms of final regret and AUC. For instance, on QAP (n=15), the Merge kernel attains a final regret of $6578.0\pm 4579.2$, significantly better than the Mallows kernel’s $8090.4\pm 4101.9$, which can be attributed to the Merge kernel’s effective capture of both local and global ordering patterns inherent in facility-location scenarios. For Cell Placement task, Merge kernel’s slightly higher regret AUC but better stability indicates a robustness advantage, especially valuable in real-world circuit design scenarios. These results confirm that our proposed kernel offers superior optimization performance and generalizes effectively to a variety of real-world permutation problems.
Ablation Study
In this section, we perform ablation experiments to systematically evaluate how the exclusion of each individual trick affects the overall effectiveness of the kernel. The ablation study results in Table 2 demonstrate that each of the three auxiliary descriptors—pairwise comparison, pattern (sliding window motif), and shift histogram—contributes to the overall performance of the Merge kernel. When all tricks are included, the model consistently achieves either the best or nearly best results in both simple final regret and regret AUC across most benchmarks, highlighting the complementary effects of these modules. Removing any single component typically leads to a performance drop, underscoring the importance of each descriptor in capturing diverse structural features of permutation problems.
We observe that removing the shift histogram module consistently results in a notable performance degradation across all tasks, as it provides a robust global positional signal, which is essential for quickly identifying globally optimal permutation structures, indicating that this component is the most impactful among the three descriptors in terms of optimization performance. Interestingly, the impact of the other two tricks varies by task. For instance, omitting the pairwise comparison module yields a lower final regret on the Cell Replacement task and the QAP task, however it also results to higher regret AUC, suggesting that local pairwise information may be a trade-off option for optimization speed and performance. In contrast, removing the pattern module leads to a noticeable decline in AUC for the Floor Planning benchmark, indicating its particular effectiveness in capturing local ordering relevant to that problem. Overall, these results confirm that the three lightweight descriptors are largely complementary, and that their joint inclusion ensures robust and superior optimization performance across a range of permutation-based tasks.
Conclusions and Limitations
In this work, we proposed a novel kernel construction framework for permutation spaces by leveraging sorting algorithms as structured comparison schemes. Within this framework, we introduced the Merge kernel—an efficient, compact, and theoretically grounded alternative to the quadratic Mallows kernel. Through extensive experiments on both synthetic and real-world benchmarks, MergeBO consistently outperformed the state-of-the-art BOPS-H algorithm in terms of optimization performance, particularly on tasks where local structure and partial orderings dominate the objective landscape. Theoretically, we showed that the Merge kernel achieves the information-theoretic lower bound on feature complexity while preserving meaningful structure from the original permutation. This insight bridges sorting theory and kernel design, opening new directions for structured optimization. The proposed framework is modular and extensible, enabling the integration of other structured comparison strategies and additional descriptors without sacrificing computational efficiency. This opens possibilities for scaling BO to larger permutation spaces and more diverse application domains.
Despite the promising results, our study still has several limitations. Our study isolates the impact of the Merge kernel by keeping BOPS-H’s neighbour-based local search and acquisition function unchanged. Consequently, most runtime is still dominated by neighbourhood exploration, so the $O(n\log n)$ feature map yields little speed-up on the modest problem sizes used here ($n\leq$30). Nor did we leverage gradient methods (e.g., L-BFGS-B) that could exploit the lower-dimensional feature space. In addition, our evaluation is confined to a single BO framework, leaving the kernel’s behaviour with alternative acquisition strategies or global search heuristics unexplored. Finally, we did not examine other fixed sorting–network algorithms such as bitonic sort because its $O(n\log n^{2})$ comparison count yields a feature vector that, for the problems with relatively small dimensions, can actually exceed the $O(n^{2})$ length of the Mallows kernel—thereby undermining our goal of dimensionality reduction. We therefore plan to (i) replace neighbour search with gradient-based acquisition optimization, (ii) applying the Merge kernel to AI tasks involving larger permutations, such as robotics planning or automated scientific discovery pipelines ($n>$100) to expose asymptotic gains, (iii) benchmark the Merge kernel within diverse BO frameworks to assess its generality, and (iv) generate and evaluate kernel function from bitonic sort algorithm on the larger permutation space.
Acknowledgments
We thank Xenophon Evangelopoulos from the University of Liverpool for drawing our attention to the lack of right‑invariance property in the Merge kernel, an insight that inspired us to devise lightweight descriptors to improve the algorithm’s performance.

### v3

@@
-Merge Kernel for Bayesian Optimization on Permutation Space
+From Sorting Algorithms to Scalable Kernels: Bayesian Optimization in High-Dimensional Permutation Spaces
 Abstract
-Bayesian Optimization (BO) algorithm is a standard tool for black-box optimization problems.
-The current state-of-the-art BO approach for permutation spaces relies on the Mallows kernel-an $\Omega(n^{2})$ representation that explicitly enumerates every pairwise comparison.
-Inspired by the close relationship between the Mallows kernel and pairwise comparison, we propose a novel framework for generating kernel functions on permutation space based on sorting algorithms.
+Bayesian Optimization (BO) is a powerful tool for black-box optimization, but its application to high-dimensional permutation spaces is severely limited by the challenge of defining scalable representations.
+The current state-of-the-art BO approach for permutation spaces relies on an exhaustive $\Omega(n^{2})$ pairwise comparison, inducing a dense representation that is impractical for large-scale permutations.
+To break this barrier, we introduce a novel framework for generating efficient permutation representations via kernel functions derived from sorting algorithms.
 Within this framework, the Mallows kernel can be viewed as a special instance derived from enumeration sort.
-Further, we introduce the Merge Kernel constructed from merge sort, which replaces the quadratic complexity with $\Theta(n\log n)$ to achieve the lowest possible complexity.
-The resulting feature vector is significantly shorter, can be computed in linearithmic time, yet still efficiently captures meaningful permutation distances.
-To boost robustness and right-invariance without sacrificing compactness, we further incorporate three lightweight, task-agnostic descriptors: (1) a shift histogram, which aggregates absolute element displacements and supplies a global misplacement signal; (2) a split-pair line, which encodes selected long-range comparisons by aligning elements across the two halves of the whole permutation; and (3) sliding-window motifs, which summarize local order patterns that influence near-neighbor objectives.
-Our empirical evaluation demonstrates that the proposed kernel consistently outperforms the state-of-the-art Mallows kernel across various permutation optimization benchmarks.
-Results confirm that the Merge Kernel provides a more compact yet more effective solution for Bayesian optimization in permutation space.
-Introduction
-As one of the most widely adopted approaches to black-box optimization, Bayesian optimization (BO) (Shahriari et al.
-2015) has found broad application in machine-learning hyper-parameter tuning (Wu et al.
-2019), financial portfolio optimization (Gonzalvez et al.
-2019), chemical‐material discovery (Luo et al.
-2025), and catalyst formulation design (Xie et al.
-2023).
+Further, we introduce the Merge Kernel , which leverages the divide-and-conquer structure of merge sort to produce a compact, $\Theta(n\log n)$ to achieve the lowest possible complexity with no information loss and effectively capture permutation structure.
+Our central thesis is that the Merge Kernel performs competitively with the Mallows kernel in low-dimensional settings, but significantly outperforms it in both optimization performance and computational efficiency as the dimension $n$ grows.
+Extensive evaluations on various permutation optimization benchmarks confirm our hypothesis, demonstrating that the Merge Kernel provides a scalable and more effective solution for Bayesian optimization in high-dimensional permutation spaces, thereby unlocking the potential for tackling previously intractable problems such as large-scale feature ordering and combinatorial neural architecture search.
+1 Introduction
+As one of the most widely adopted approaches to black-box optimization, Bayesian optimization (BO) Shahriari et al.
+(2015) has found broad application in machine-learning hyper-parameter tuning Wu et al.
+(2019), financial portfolio optimization Gonzalvez et al.
+(2019), chemical and material discovery Luo et al.
+(2025), and catalyst formulation design Xie et al.
+(2023).
 BO employs probabilistic surrogate models—most commonly Gaussian processes (GPs)—to approximate the unknown objective and uses an acquisition function to balance exploration and exploitation, thereby approaching the global optimum with the fewest possible evaluations.
-Most research on BO to date has concentrated on continuous (Greenhill et al.
-2020) and categorical design spaces (Garrido-Merchán and Hernández-Lobato 2020; Nguyen et al.
-2020; Bartoli et al.
-2025), whereas applications to permutation spaces remain comparatively underexplored.
-This gap is striking given the ubiquity of permutation optimization in both theory and practice: canonical examples include the travelling-salesperson problem (TSP), the sequencing of operations in automated experimental pipelines (Guidi, Seeberger, and Gilmore 2020), and the sequential order-of-addition experiments (Lin and Rios 2025).
-In addition, permutation optimization frequently arises in diverse AI applications, including scheduling tasks in robotic planning (Alatartsev, Stellmacher, and Ortmeier 2015), optimizing experimental sequences (Blau et al.
-2022), and other sequential decision-making problems (Sun and Giles 2001; Wen et al.
-2023).
+Most research on BO to date has concentrated on continuous Greenhill et al.
+(2020) and categorical design spaces Garrido-Merchán & Hernández-Lobato (2020); Nguyen et al.
+(2020); Bartoli et al.
+(2025), whereas applications to permutation spaces remain comparatively underexplored.
+This gap is striking given the ubiquity of permutation optimization in both theory and practice: canonical examples include the traveling-salesperson problem (TSP), the sequencing of operations in automated experimental pipelines Guidi et al.
+(2020), and the sequential order-of-addition experiments Lin & Rios (2025).
+In addition, permutation optimization frequently arises in diverse AI applications, including scheduling tasks in robotic planning Alatartsev et al.
+(2015), optimizing experimental sequences Blau et al.
+(2022), and other sequential decision-making problems Sun & Giles (2001); Wen et al.
+(2023).
 Therefore, extending Bayesian optimization to permutation spaces holds significant theoretical and practical value for artificial intelligence as well as broader academic and industrial applications.
-Successfully deploying BO in permutation spaces hinges on equipping the GP surrogate with a kernel that faithfully quantifies the similarity between two permutations.
-Existing approaches have evolved along two principal lines: (1) General-purpose discrete BO frameworks—most notably COMBO (Oh et al.
-2019).
-COMBO combines a graph-Laplacian kernel with local search to accommodate heterogeneous discrete variables and thus enjoys broad applicability.
-However, it relies on manually crafted adjacency graphs or hash encodings, which struggle to capture the comparison sequences and cyclic shifts unique to permutations; consequently, its efficiency degrades on tasks that involve frequent high-order swaps or mixed local/global reorderings.
-(2) Permutation-specific kernels—chief among them the Mallows kernel in BOPS(Deshwal et al.
-2022).
-This kernel is built on the Kendall–$\tau$ distance, representing a permutation via the inversion counts generated by enumeration sort, yielding a feature dimension of $O(n^{2})$.
-The dimension therefore scales quadratically with $n$, and because $2^{n^{2}}\!\gg n!$, the vast majority of features fail to correspond to any valid permutation, resulting in both statistical and computational redundancy.
+[…]
 To address the limitations of existing approaches, we propose a sorting-algorithm–driven kernel-design framework for permutations and instantiate it with the Merge kernel, which reduces the feature dimension to $O(n\log n)$—the information-theoretic lower bound for encoding a permutation.
@@
 Choosing an algorithm with a deterministic comparison tree—such as merge sort or bitonic sort—thus produces a representation that is both fixed in length and highly compact.
-To further enhance robustness and expressiveness, we augment the Merge kernel with three lightweight descriptors:
-  • Shift Histogram – captures invariance under cyclic shifts;
-  • Split-Pair Line – emphasises the relative order of elements across partition boundaries;
-  • Sliding-Window Motif – extracts local sorting patterns, increasing sensitivity to local perturbations.
-This design attains the shortest possible encoding while simultaneously eliminating the high-dimensional redundancy of the Mallows kernel and the structural blind spots of COMBO, thereby providing an efficient and accurate similarity measure for Bayesian optimization over permutation spaces.
 Contributions. Our work makes three principal contributions:
@@
 Within this view, the classic Mallows kernel is recovered as the special case obtained when the framework is instantiated with enumeration sort.
-  • Merge kernel.
-Applying the framework to merge sort yields Merge Kernel, whose $O(n\log n)$ construction matches the information-theoretic lower bound on comparison complexity.
-We further enrich it with three lightweight descriptors to boost predictive power at minimal cost.
+  • Merge kernel. Applying the framework to merge sort yields Merge Kernel, whose $O(n\log n)$ construction matches the information-theoretic lower bound on comparison complexity.
   • Comprehensive evaluation.
-We assess the effectiveness of our kernels on diverse synthetic and real-world benchmarks, consistently outperforming the state-of-the-art Mallows kernel in both accuracy and efficiency.
+We assess the effectiveness of our kernels on diverse synthetic and real-world benchmarks.
+Results on low-dimensional benchmarks show competitive performance against the state-of-the-art Mallows kernel, while it significantly outperforms the Mallows kernel on high-dimensional benchmarks.
 Our results demonstrate that the Merge kernel provides a practical and efficient tool for permutation optimization, significantly enhancing BO’s applicability to diverse AI scenarios.
-Background and Related Works
-Permutation Optimization
-Here we describe the problem formulation of permutation optimization with a fixed length $n\in\mathcal{N}$.
-Let $[n]=\{1,2,...,n\}$, a permutation is a function $\pi:[n]\longrightarrow[n]$ such that $\pi$ is bijective.
+2 Background and Related Works
+2.1 Permutation Optimization
+Here we describe the problem formulation of permutation optimization with a fixed length $n\in\mathbb{N}$.
+Let $n]=\{1,2,...,n\}$, a permutation is a function $\pi:[n]\longrightarrow[n]$ such that $\pi$ is bijective.
 The set of all permutations of $[n]$ is the symmetric group
 $\mathcal{S}_{n}\;=\;\bigl\{\pi\mid\pi:[n]\longrightarrow[n]\text{ is bijective}\bigr\}.$
-We are given a costly-to-evaluate, possibly noisy black-box function
-$f:\;\mathcal{S}_{n}\longrightarrow\mathbb{R},$
-which assigns a real-valued quality (e.g., cost, loss, reward) to every permutation $\pi$. Hence, the optimization problem can be formulated as
+We are given a costly-to-evaluate, possibly noisy black-box function $f:\;\mathcal{S}_{n}\longrightarrow\mathbb{R}$, which assigns a real-valued quality (e.g., cost, loss, reward) to every permutation $\pi$.
+Hence, the optimization problem can be formulated as
 $\pi^{*}\;=\;\text{arg min}_{\pi\;\in\;\mathcal{F}}f(\pi)$
@@
 In this study we only consider the unconstrained case, therefore we have $\mathcal{F}\;=\;\mathcal{S}_{n}$; in practice $\mathcal{F}$ may exclude permutations violating domain rules.
-Bayesian Optimization
-Bayesian optimization (BO) (Shahriari et al.
-2015) is an optimization algorithm for black-box objective functions that no closed-form expression or gradient information is available and whose evaluation is often an expensive physical or computational experiment.
-The algorithm first fits the observed data with a surrogate model, most commonly a Gaussian process (GP) (Williams and Rasmussen 2006), and then employs an acquisition function to select the next query point, balancing exploration of uncertain regions against exploitation of promising areas identified by the surrogate.
+2.2 Bayesian Optimization
+Bayesian optimization (BO) Shahriari et al.
+(2015) is an optimization algorithm for black-box objective functions that no closed-form expression or gradient information is available and whose evaluation is often an expensive physical or computational experiment.
+The algorithm first fits the observed data with a surrogate model, most commonly a Gaussian process (GP) Williams & Rasmussen (2006), and then employs an acquisition function to select the next query point, balancing exploration of uncertain regions against exploitation of promising areas identified by the surrogate.
 The kernel $K(\mathbf{x},\mathbf{x}^{\prime})$ is the central design lever in a Gaussian-process surrogate: it defines the similarity metric between inputs, thereby specifying the prior smoothness assumptions and, through GP inference, the posterior mean and uncertainty.
@@
 Whereas Euclidean spaces typically rely on Gaussian (RBF) kernels, discrete structures—and permutations in particular—require bespoke constructions that faithfully encode ordering relationships.
-In the next subsection we revisit the Mallows kernel on the symmetric group $\mathcal{S}_{n}$, outline its computational and statistical shortcomings, and use these observations to motivate our own kernel design.
-Mallows Kernel for Permutation Space
-BOPS-H (Deshwal et al.
-2022) is the current state-of-the-art BO algorithm for permutation optimization, which propose to employ Mallows kernel (Jiao and Vert 2015) on the symmetric group $\mathcal{S}_{n}$ in a similar manner to the RBF kernel on the Euclidean space.
+For example, Oh et al.
+(2022) uses the Position kernel Zaefferer et al.
+(2014) and L-ensemble with Acquisition Weights to extend BO on permutation space to a batched BO scheme.
+While significant progress has been made in scaling BO to high-dimensional continuous domains (e.g., Eriksson et al.
+(2019)Wang et al.
+(2016)), extending BO to large-scale structured discrete domains like permutations presents a distinct set of challenges centered on kernel design, which is the primary focus of this work.
+2.3 Mallows Kernel for Permutation Space
+BOPS-H Deshwal et al.
+(2022) is the current state-of-the-art BO algorithm for permutation optimization, which proposes to employ Mallows kernel Jiao & Vert (2015) on the symmetric group $\mathcal{S}_{n}$ in a similar manner to the RBF kernel on the Euclidean space.
 The Mallows kernel $K_{mal}(\pi,\pi^{\prime})$ for the permutation pair ($\pi$, $\pi^{\prime}$) is defined as the exponential negative of the number of discordant pairs $n_{d}(\pi,\pi^{\prime})$ between $\pi$ and $\pi^{\prime}$:
 $K_{Mal}(\pi,\pi^{\prime})=\exp(-ld(\pi,\pi^{\prime}))$ (1)
-where $l\geq 0$ is the length-scale parameter of the Mallows kernel, and $d(\pi,\pi^{\prime})$ is the Kendall-$\tau$ distance (Kendall 1938) which counts the number of pairs of elements ordered oppositely by $\pi$ and $\pi^{\prime}$:
+where $l\geq 0$ is the length-scale parameter of the Mallows kernel, and $d(\pi,\pi^{\prime})$ is the Kendall-$\tau$ distance Kendall (1938) which counts the number of pairs of elements ordered oppositely by $\pi$ and $\pi^{\prime}$:
 $d\left(\pi,\pi^{\prime}\right)=\sum_{i<j}$ $\left[1_{\pi(i)>\pi(j)}1_{\pi^{\prime}(i)<\pi^{\prime}(j)}\right$. (2)
@@
 Two pairs are discordant among the six unordered pairs: $(1,2),(3,4)$, hence $d(\pi,\pi^{\prime})=2$ and $K_{Mal}(\pi,\pi^{\prime})=\exp(-2l)$.
-Algorithm
-Generating Kernel based on Sorting Algorithm
+3 Merge Kernel: Generating Kernels from sorting algorithms
+Algorithm 1 Merge feature mapping $\Phi_{Mer}(\pi)$
+Input: Permutation $\pi$ of length $n$
+Output: Feature vector $\Phi_{Mer}(\pi)$
+if length($\pi$) == 1 then
+  return []
+end if
+if length($\pi$) == 2 then
+  return [1] if $\pi[0]>\pi$[1] else [0]
+end if
+ Let $mid=\lfloor\frac{n}{2}\rfloor$
+ Let $V_{Left},V_{Right}=\Phi_{Mer}(\pi[:mid]),\Phi_{Mer}(\pi[mid:])$
+ Let $\hat{\pi}_{l},\hat{\pi}_{r}=\text{sorted}(\pi[:mid]),\text{sorted}(\pi[mid:])$
+ Let $V_{Merge}=[]$, $i=j=0$
+while $i<\text{length}(\hat{\pi}_{l})$ and $j<\text{length}(\hat{\pi}_{r})$ do
+  if $\hat{\pi}_{l}[i]>\hat{\pi}_{r}[j]$ then
+   $V_{Merge}$.append(1)
+   $j+=1$
+  end if
+  if $\hat{\pi}_{l}[i]<\hat{\pi}_{r}[j]$ then
+   $V_{Merge}$.append(0)
+   $i+=1$
+  end if
+end while
+return $V_{Left}+V_{Right}+V_{Merge}$
 In the previous section, we have shown that the core of the Mallows kernel is the pairwise comparison of all elements. Equivalently, it maps a permutation $\pi$ to a feature vector
-$\Phi_{Mal}(\pi)\;\in\;\{-1,1\}^{\binom{n}{2}}$
-where each coordinate corresponds to the comparison of a pair of elements: -1 if they are in ascending order, and +1 otherwise. We then have
-$\begin{split}K_{Mal}(\pi,\sigma)&=\exp\!\Bigl{(}-\frac{\|\Phi_{Mal}(\pi)-\Phi_{Mal}(\sigma)\|^{2}}{2\ell^{2}}\Bigr{)}\\ &=\;k_{\mathrm{RBF}}\bigl{(}\Phi_{Mal}(\pi),\,\Phi_{Mal}(\sigma)\bigr{)}.\end{split}$ (3)
-Since the Gaussian RBF kernel $k_{\mathrm{RBF}}$ is strictly positive definite on $\mathbb{R}^{d}$ and thus satisfies Mercer’s condition (Mercer 1909), and because positive definiteness is preserved under composition with any deterministic mapping $\Phi$, it follows that $K(\pi,\sigma)$ constructed from $\Phi$ also satisfies Mercer’s condition and is therefore a valid kernel function.
+$\Phi_{Mal}(\pi)\;\in\;\{0,1\}^{\binom{n}{2}}$
+where each coordinate corresponds to the comparison of a pair of elements: 0 if they are in ascending order, and 1 otherwise. We then have
+$\begin{split}K_{Mal}(\pi,\pi^{\prime})&=\exp\!\Bigl(-\frac{\|\Phi_{Mal}(\pi)-\Phi_{Mal}(\pi^{\prime})\|^{2}}{2\ell^{2}}\Bigr)\\ &=\;K_{\mathrm{RBF}}\bigl(\Phi_{Mal}(\pi),\,\Phi_{Mal}(\pi^{\prime})\bigr).\end{split}$ (3)
+under the reparameterization $l=\frac{1}{2\ell^{2}}$ for the length-scale parameter $\ell$ in the Gaussian RBF kernel.
+Since the RBF kernel $K_{\mathrm{RBF}}$ is strictly positive definite on $\mathbb{R}^{d}$ and thus satisfies Mercer’s condition Mercer (1909), and because positive definiteness is preserved under composition with any deterministic mapping $\Phi$, it follows that $K(\pi,\pi^{\prime})$ constructed from $\Phi$ also satisfies Mercer’s condition and is therefore a valid kernel function.
 Thus, other pairwise comparison methods can also be used to construct analogous feature vectors, which—when combined with an RBF kernel—yield valid kernel functions.
@@
 This is because the mapping from permutations to feature space must yield feature vectors of fixed length; otherwise, feature vectors of differing lengths would not be compatible with the RBF kernel.
-Hence, only sorting algorithms that have a fixed comparison path and a constant number of comparisons across all inputs can generate valid feature mappings.
+Hence, only sorting algorithms that have a fixed comparison path and a constant number of comparisons across all inputs can generate valid feature mappings, e.g., a fixed sorting network with a predetermined comparator sequence Batcher (1968).
 In fact, most sorting algorithms commonly known to have complexity of order $O(n\log n)$—such as quicksort, heapsort, and standard merge sort—do not satisfy the requirement of constant comparison counts.
-Nevertheless, merge sort is an exception: although typically merge sort ceases comparisons once all elements from one subsequence have been merged, redundant comparisons can be artificially introduced during the merge procedure.
+Nevertheless, merge sort could be an exception: although typically merge sort ceases comparisons once all elements from one subsequence have been merged, redundant comparisons can be artificially introduced during the merge procedure, resulting in a fixed number of comparisons L+R1L+R-1, where $L$ and $R$ are the length of the two subsequences.
 This ensures that both the comparison path and the number of comparisons remain identical across different permutations, thus establishing a fixed comparison path.
 Consequently, we select merge sort as the feature mapping $\Phi_{Mer}(\pi)$ for constructing permutation-space kernel function $K_{Mer}$, following Equation 3:
-$\begin{split}K_{Mer}(\pi,\sigma)&=\exp\!\Bigl{(}-\frac{\|\Phi_{Mer}(\pi)-\Phi_{Mer}(\sigma)\|^{2}}{2\ell^{2}}\Bigr{)}\\ &=\;k_{\mathrm{RBF}}\bigl{(}\Phi_{Mer}(\pi),\,\Phi_{Mer}(\sigma)\bigr{)}.\end{split}$ (4)
-The element-pair comparison mapping $\Phi_{Mer}(\pi)$ is shown in Algorithm 1.
-Algorithm 1 Merge feature mapping $\Phi_{Mer}(\pi)$
-Input: Permutation $\pi$ of length $n$
-Output: Feature vector $\Phi_{Mer}(\pi)$
-  if length($\pi$) == 1 then
-     return []
-  end if
-  if length($\pi$) == 2 then
-     return [1] if $\pi[0]>\pi$[1]
-     return [-1] if $\pi[0]<\pi$[1]
-  end if
-  Let $mid=\lfloor\frac{n}{2}\rfloor$
-  Let $V_{Left}=\Phi_{Mer}(\pi[:mid])$
-  Let $V_{Right}=\Phi_{Mer}(\pi[mid:])$
-  Let $\hat{\pi}_{l}=\text{sorted}(\pi[:mid])$
-  Let $\hat{\pi}_{r}=\text{sorted}(\pi[mid:])$
-  Let $V_{Merge}=[]$
-  Let $i=j=0$
-  while $i<\text{length}(\hat{\pi}_{l})$ and $j<\text{length}(\hat{\pi}_{r})$ do
-     if $\hat{\pi}_{l}[i]>\hat{\pi}_{r}[j]$ then
-        $V_{Merge}$.append(1)
-        $j+=1$
-     end if
-     if $\hat{\pi}_{l}[i]<\hat{\pi}_{r}[j]$ then
-        $V_{Merge}$.append(-1)
-        $i+=1$
-     end if
-  end while
-  return $V_{Left}+V_{Right}+V_{Merge}$
+$\begin{split}K_{Mer}(\pi,\pi^{\prime})&=\exp\!\Bigl(-\frac{\|\Phi_{Mer}(\pi)-\Phi_{Mer}(\pi^{\prime})\|^{2}}{2\ell^{2}}\Bigr)\\ &=\;K_{\mathrm{RBF}}\bigl(\Phi_{Mer}(\pi),\,\Phi_{Mer}(\pi^{\prime})\bigr).\end{split}$ (4)
+The element-pair comparison mapping $\Phi_{Mer}(\pi)$ is shown in Algorithm 1. Here we present an example for permutation [1,4,3,2] to show the Merging process.
+Example: Feature Mapping for $\pi=(1,4,3,2)$
+  1. 1.
+    Initial Split: $\pi$ splits into $L_{0}=(1,4)$ and $R_{0}=(3,2)$.
+  2. 2.
+    Recurse on $L_{0}=(1,4)$: Merging sorted lists ‘[1]’ and ‘[4]’ yields feature vector $V_{L}=[0]$.
+  3. 3.
+    Recurse on $R_{0}=(3,2)$: Merging sorted lists ‘[2]’ and ‘[3]’ yields feature vector $V_{R}=[1]$.
+  4. 4.
+    Final Merge: Merge sorted lists ‘(1,4)’ and ‘(2,3)’. The fixed comparison path generates the merge vector $V_{Merge}=[0(1<2),1(4>2),1(4>3),1(\text{left padding})]$.
+  5. 5.
+    Concatenate: The final feature vector is $\Phi_{Mer}(\pi)=V_{L}\oplus V_{R}\oplus V_{Merge}=[0]\oplus[1]\oplus[0,1,1,1]=[0,1,0,1,1,1]$.
 We have established that merge sort, with a specially designed fixed-comparison procedure, is uniquely capable of constructing valid kernel functions among common sorting algorithms with a complexity of $\Omega(n\log n)$.
@@
 On the other hand, consider the permutation space consisting of all $n!$ possible permutations of length $n$.
-From an information-theoretic viewpoint, encoding all $n!$ permutations without loss using a binary feature vector composed solely of ${-1,1}$ requires a minimum vector length of $\log_{2}(n!)$.
-Applying Stirling’s approximation (Donald et al.
-1999), we have:
-$\log_{2}(n!)=n\log_{2}n-nlog_{2}e+O(\log n)=\Omega(n\log n)$
+From an information-theoretic viewpoint, encoding all $n!$ permutations without loss using a binary feature vector composed solely of 0,1{0,1} requires a minimum vector length of $\log_{2}(n!)$.
+Applying Stirling’s approximation Donald et al.
+(1999), we have: $\log_{2}(n!)=n\log_{2}n-nlog_{2}e+O(\log n)=\Omega(n\log n)$.
 Consequently, the feature vector length of the merge-sort-based kernel (Merge Kernel) reaches this theoretical lower bound for lossless permutation encoding.
-Adding Invariance to Merge Kernel
+[…]
 It is worth noting that, when constructing the Merge kernel via merge sort, we have not required the feature mapping $\Phi$ to possess any group invariance property, such as right-invariance.
 Traditionally, a permutation-distance measure should be invariant under right multiplication, meaning that applying an identical right-multiplication operation to two permutations should not alter the distance between them.
-However, only sorting algorithms with complexity $O(n^{2})$ can yield fully right-invariant kernel functions, since such invariance necessitates exhaustive pairwise comparisons among all $\frac{n(n-1)}{2}$ pairs of elements—an impossibility for more efficient sorting algorithms like merge sort.
+However, only sorting algorithms with complexity $O(n^{2})$ can yield fully right-invariant kernel functions with no information loss (otherwise, a simple Spearman’s footrule Diaconis & Graham (1977) with $O(n)$ complexity can hold right-invariance as well), since such invariance necessitates exhaustive pairwise comparisons among all $\frac{n(n-1)}{2}$ pairs of elements—an impossibility for more efficient sorting algorithms like merge sort.
 Consequently, although the Merge kernel achieves better computational efficiency through a more compact encoding, it sacrifices a certain degree of performance due to the loss of right-invariance.
-To address this issue, we can naturally consider trading off some compactness by slightly increasing the encoding length—while preserving the overall $\Omega(n\log n)$ complexity—to enhance right-invariance and thus compensate for the associated performance loss.
-We introduce three additional encoding strategies, whose resulting feature vectors are concatenated to the original Merge feature mapping to further improve the kernel’s performance.
-Algorithm 2 Sliding-window Motif Feature Mapping $\Phi_{Sli}(\pi)$
-Input: permutation $\pi$ of length $n$; window length $w$
-Output: sliding-window motif vector $\Phi_{Sli}(\pi)$
-[…]
-  build dictionary $\text{index}[p]\leftarrow$ rank of pattern $p$ in patterns
-  $H\leftarrow\mathbf{0}_{\,w!}$initialise histogram}
-  for $i\leftarrow 0$ to $n-1$ do
-     $\text{win}\leftarrow\bigl{(}\pi_{(i+j)\bmod n}\bigr{)}_{j=0}^{w-1}$circular slice}
-     $p\leftarrow\textsc{Relative Order}(\text{win})$
-     $idx\leftarrow\text{index}[p]$
-     $H[idx]\leftarrow H[idx]+1$
-  end for
-  return $H\;\;(=\Phi_{Sli}(\pi))$
-Algorithm 3 Relative Order$(v)$
-Input: sequence $v=(v_{0},\dots,v_{w-1})$
-Output: permutation $p$ giving the relative order of $v$
-[…]
-  $p\leftarrow$ array of length $w$
-[…]
-     $p[i]\leftarrow r$
-  end for
-  return tuple$(p)$
-Algorithm 4 Shift-histogram feature mapping $\Phi_{Shi}(\pi)$
-Input: Permutation $\pi$ of length $n$, maximum shift $s_{\max}$
-Output: Shift-histogram feature vector $\Phi_{Sli}(\pi)$
-  $n\leftarrow\text{length}(\pi)$
-  for $i\leftarrow 0$ to $n-1$ do
-     $\text{pos}[\pi[i]]\leftarrow i$
-  end for
-  for $i\leftarrow 0$ to $n-1$ do
-     $\Delta_{i}\leftarrow\text{pos}[i]-i$
-     if $\Delta_{i}>s_{\max}$ then
-        $\Delta_{i}\leftarrow s_{\max}$
-     else if $\Delta_{i}<-s_{\max}$ then
-        $\Delta_{i}\leftarrow-s_{\max}$
-     end if
-  end for
-[…]
-  for $i\leftarrow 0$ to $n-1$ do
-     $H[\Delta_{i}+s_{\max}]\leftarrow H[\Delta_{i}+s_{\max}]+1$
-  end for
-  return $H$
-Middle-split cross-column pair comparison
-Although merge sort efficiently encodes comparisons within local subsets, it does not explicitly cover cross-partition pairs at each recursive split, specifically the first split—the comparison vector $V_{Merge}$ will be computed in the last recursion after left half and right half sorted.
-To remedy this and capture long-range interactions, we propose to add the strict pairwise comparisons $\Phi_{Mid}(\pi)$ between elements sharing the same offset in their respective halves after first split.
-For example, $\Phi_{Mid}(\pi)$ will have 3 digits when $\pi$ has length of 6: comparison between $(\pi_{0},\pi_{3}),(\pi_{1},\pi_{4}),(\pi_{2},\pi_{5})$.
-To keep the representation compact, we retain only the cross-partition comparisons produced by the outermost split.
-At this level the two halves are merged only in the final recursion step, so their relative order is maximally perturbed and the partition gap is the largest, encapsulating the richest set of inter-partition comparisons while requiring only $\Omega(n)$ encoding length.
-Because the resulting procedure is straightforward, we omit its pseudocode here.
-Sliding-window motif
-To capture local ordering patterns that heavily influence near-neighbor-dependent objectives and add cyclic-rotation invariance, we introduce sliding-window motifs.
-For a small number of window length $k$, we enumerate all possible local ordering patterns (motifs) and count their occurrences within the permutation.
-With three-element windows, there are exactly $k!$ distinct relative-order patterns.
-We store their frequencies as a compact local-order texture descriptor, further enriching the kernel with information on short-range permutation structures.
-Because the sliding window wraps around and fills from the beginning of the permutation once it reaches the end, this scheme explicitly endows the features with cyclic-rotation invariance.
-The feature mapping algorithm $\Phi_{Sli}(\pi)$ is shown in Algorithm 2 and 3.
-Shift-histogram
-To incorporate another global misplacement measure that remains invariant under cyclic right-multiplication, we introduce the shift histogram $\Phi_{Shi}$.
-For each permutation $\pi$, we first calculate the absolute displacement of each element relative to its position in the identity permutation, defined formally as:
-$d_{i}=\pi^{-1}(i)-i,\forall i\in\{0,1,...,n-1\}$
-The displacements are then binned into a fixed-size histogram, typically truncated at some maximum absolute shift (e.g.,±$\pm 5$), resulting in a compact, stable global descriptor of permutation misplacement.
-The feature mapping algorithm $\Phi_{Shi}(\pi)$ is shown in Algorithm 4.
-Eventually, the final Merge kernel uses a concatenated feature mapping strategy contains all elements above:
-$\Phi_{Merge}(\pi)=concat(\Phi_{Mer}(\pi),\Phi_{Mid}(\pi),\Phi_{Sli}(\pi),\Phi_{Shi}(\pi))$
-Intuitively, these lightweight descriptors enhance the kernel’s ability to capture global and local permutation structures, improving both optimization stability and robustness.
-Experiments
-Figure 1: Results comparing Mallows kernel (BOPS-H) and Merge kernel (MergeBO) on the current regret value (difference between best-so-far and optimal value) vs. number of iteration.
-Simple final regret Best so far AUC
-Merge Mallows Merge Mallows
-$\text{TSP}_{n=15}$ 0.012$\pm$0.038 0.012$\pm$0.038 400.9$\pm$115.0 428.2$\pm$121.9
-$\text{QAP}_{n=15}$ 6578.0$\pm$4579.2 8090.4$\pm$4101.9 2493028.7$\pm$690320.7 2750160.0$\pm$744154.9
-$\text{FloorPlanning}_{n=30}$ 23.3$\pm$11.5 30.1$\pm$12.8 8140.8$\pm$2386.8 8665.7$\pm$2638.8
-$\text{CellReplacement}_{n=30}$ 6.1$\pm$2.1 6.1$\pm$3.0 5371.0$\pm$792.7 5350.5$\pm$910.1
-Table 1: Performance comparison between MergeBO and BOPS-H (Mallows kernel), better performance shown in bold font
-Simple final regret All tricks No pairwise No pattern No shift
-$\text{TSP}_{n=15}$ 0.012$\pm$0.038 0.025$\pm$0.051 0.038$\pm$0.059 0.019$\pm$0.046
-$\text{QAP}_{n=15}$ 6578.0$\pm$4579.2 6330.0 $\pm$3132.2 6428.2$\pm$3777.5 8280.6$\pm$4536.0
-$\text{FloorPlanning}_{n=30}$ 23.3$\pm$11.5 25.4$\pm$9.4 26.9$\pm$11.5 26.2$\pm$10.5
-$\text{CellReplacement}_{n=30}$ 6.1$\pm$2.1 5.1$\pm$2.1 6.1$\pm$2.4 6.3$\pm$3.7
-Best so far AUC All tricks No pairwise No pattern No shift
-$\text{TSP}_{n=15}$ 400.9$\pm$115.0 418.7$\pm$125.3 471.6$\pm$167.5 410.8$\pm$107.3
-$\text{QAP}_{n=15}$ 2493028.7$\pm$690320.7 2588481.3$\pm$638766.7 2580381.5$\pm$838626.7 2975522.9$\pm$743481.9
-$\text{FloorPlanning}_{n=30}$ 8140.8$\pm$2386.8 8311.2$\pm$2202.0 7940.6$\pm$1908.8 8379.9$\pm$2237.7
-$\text{CellReplacement}_{n=30}$ 5371.0$\pm$792.7 5553.1$\pm$1096.4 5390.3$\pm$952.2 5590.6$\pm$1040.5
-Table 2: Ablation experiment results with each trick dropped, better performance shown in bold font
-In this section, we demonstrate the benchmarks, the experimental settings, the ablation studies as well as the results and discussion.
-Benchmarks and Experiment Settings
-Because our method, MergeBO, is a direct generalisation of the state-of-the-art Mallows kernel in BOPS-H algorithm(Deshwal et al.
-2022), we restrict the empirical evaluation to that baseline and adopt the identical benchmark suite and experimental protocol introduced in (Deshwal et al.
-2022).
-This isolates the effect of replacing the Mallows kernel with our merge-based variant and eliminates confounding factors due to data splits or hyper-parameter choices.
+4 Experiments
+4.1 Benchmarks and Experiment Settings
+Our empirical evaluation focuses on the state-of-the-art BOPS-H algorithm Deshwal et al.
+(2022) as the sole baseline.
+This choice is twofold: first, BOPS-H was shown to substantially outperform other permutation-specific methods like COMBO Oh et al.
+(2019).
+Second, our core objective is a principled comparison between our Merge Kernel and the Mallows Kernel—both natively designed for permutations.
+We therefore exclude methods for high-dimensional continuous BO (e.g., TuRBO, REMBO) to ensure a direct evaluation of representation power without confounding factors from domain adaptation.
+To comprehensively test our scalability hypothesis, we evaluate both methods on a suite of low- and high-dimensional benchmarks.
+4.1.1 Low-dimensional Benchmarks
+The study by Deshwal et al.
+(2022) exclusively considers problems with dimensions of 30 or less.
+Accordingly, we adopt their experimental settings to form our suite of low-dimensional benchmarks.
 The BOPS-H algorithm follows and modifies the local-search strategy used in COMBO, examining only the set of neighbouring permutations of the current incumbent to restrict the combinatorial search space, we therefore adopt the same procedure in our experiments.
-To keep the extra features introduced by the tricks compact and prevent them from overshadowing the Merge kernel, we set the sliding‑window length to 4 and capped the shift‑histogram at a maximum shift of 5 for all four tasks.
-GPyTorch (Gardner et al.
-2018) and BoTorch (Balandat et al.
-2020) libraries are used to implement both alogithms.
-We evaluate our method on the same two synthetic benchmarks and the same two real-world applications in (Deshwal et al.
-2022).
-Detailed information for all benchmarks are listed below:
-(1)Quadratic Assignment (QAP).
+GPyTorch Gardner et al.
+(2018) and BoTorch Balandat et al.
+(2020) libraries are used to implement both algorithms.
+Expected Improvement acquisition function is used for all the experiments, and 10 restarts are used for local search based EI optimization for BOPS-H and MergeBO.
+Each benchmark is evaluated with 20 independent trials, each consisting of 50 iterations.
+The random seed for each trial is set to its trial index.
+We evaluate our method on the same two synthetic benchmarks and the same two real-world applications in Deshwal et al. (2022). Detailed information for all benchmarks are listed below:
+(1)Quadratic Assignment (QAP)n=15.
 This is a classic facility-location problem: assign $n$ facilities to $n$ locations so that flow costs and distances align optimally.
-We use the 15-city instances from QAPLIB (Burkard, Karisch, and Rendl 1997), each with cost matrix $A$ and distance matrix $B$, and minimise $\operatorname{Tr}(APBP^{\mathsf{T}})$ over permutation matrices $P$.
-(2)Travelling Salesman (TSP).
+We use the 15-city instances from QAPLIB Burkard et al.
+(1997), each with cost matrix $A$ and distance matrix $B$, and minimise $\operatorname{Tr}(APBP^{\mathsf{T}})$ over permutation matrices $P$.
+(2)Travelling Salesman (TSP)n=15.
 The TSP seeks the shortest Hamiltonian cycle through a set of cities and is a standard benchmark for route-planning.
-Our instances are ten-city PCB drill tours from TSPLIB (Reinelt 1995); the score is the total travel time to visit all holes exactly once and return.
-(3)Floor Planning (FP).
+Our instances are 15 city PCB drill tours from TSPLIB Reinelt (1995); the score is the total travel time to visit all holes exactly once and return.
+(3)Floor Planning (FP)n=30.
 Floor planning is an NP-hard VLSI layout task that packs rectangular modules on a chip while minimising area and manufacturing cost.
-We evaluate two ten-block variants (FP-1 and FP-2); each permutation defines a block placement whose cost we minimise.
-(4)Cell Placement (CP).
-Cell placement arranges logic cells on a row to reduce wire-length and hence circuit delay.We consider ten equal-height cells with a fixed net-list; the objective is the total Manhattan wire-length induced by a permutation of cell positions.
-Because the publicly available implementation of the Mallows kernel (https://github.com/aryandeshwal/BOPS) does not provide the interface required to run the Rodinia’s heterogeneous many-core benchmark (Che et al.
-2009), we did not perform experiments on this benchmark.
-Expected Improvement acquisition function is used for all the experiments, and 10 restarts are used for local search based EI optimization for BOPS-H and MergeBO.
+We evaluate two 15-block variants (FP-1 and FP-2); each permutation defines a block placement whose cost we minimise.
+(4)Cell Placement (CP)n=30.
+Cell placement arranges logic cells on a row to reduce wire-length and hence circuit delay.
+We consider 30 equal-height cells with a fixed net-list; the objective is the total Manhattan wire-length induced by a permutation of cell positions.
+Because the publicly available implementation of the Mallows kernel (https://github.com/aryandeshwal/BOPS) does not provide the interface required to run the Rodinia’s heterogeneous many-core benchmark Che et al.
+(2009), we did not perform experiments on this benchmark.
+We note a numerical discrepancy between our replicated results and those reported in the original paper, which we attribute to subtle implementation details not specified in the publication, such as problem instance choices.
+However, we emphasize that within our experimental framework, both the Merge Kernel and the Mallows Kernel were evaluated under identical conditions, ensuring a fair and controlled comparison of their relative performance.
+4.1.2 High-dimensional Benchmarks: Traveling Thief Problems
+We introduce traveling thief problems (TTP) Bonyadi et al.
+(2013) as high-dimensional benchmarks, which is a defined as a combination of TSP and knapsack problem: a thief must determine a tour through $n$ cities with distance matrix $D=\{d_{ij}\}$ while simultaneously selecting items of varying weights $w_{k}$ and values $p_{k}$ to maximize profit without exceeding a knapsack capacity $W$.
+This structure defines a complex, hybrid search space, combining an $n$-dimensional permutation space for the city tour with a $\{0,1\}^{m}$ discrete space for item selection.
+Despite its typical application in evaluating white-box or heuristic algorithms Polyakovskiy et al.
+(2014); Gupta et al.
+(2015), we adapt the TTP as a true black-box benchmark, providing no structural information to the optimizer.
+Following the implementation in Polyakovskiy et al.
+(2014), we create three distinct instances based on a $n=280$-city problem (a 280-dimensional permutation space).
+These instances feature a large number of items with varying properties: (1) $m=279$ items with uncorrelated weights; (2) $m=837$ items with bounded strong correlation in the weights; and (3) $m=837$ items with uncorrelated weights.
+These benchmarks provide a strenuous test for our proposed MergeBO and the baseline BOPS-H.
+Notice that the TTP is a hybrid space problem, we modified both MergeBO and BOPS-H by multiplying an RBF kernel on the $\{0,1\}^{m}$ discrete space:
+$K_{TTP}((\pi,\sigma),(\pi^{\prime},\sigma^{\prime}))=K(\pi,\pi^{\prime})K_{RBF}(\sigma,\sigma^{\prime})$ (5)
+Where $\sigma$ is the item picking strategy.
+The neighbouring permutation search method for BOPS-H is computationally infeasible on such a vast space.
+Therefore, we adopt a continuous relaxation approach, treating the entire set of optimization variables $(\pi,\sigma)$ as a continuous vector for gradient-based optimization, and the result is subsequently projected back to the nearest permutation and binary vectors.
+All experiments use the EI acquisition function.
+Each benchmark is evaluated with 50 independent trials, each consisting of 55 iterations (5 iterations of random initialization and 50 iterations of optimization).
+The random seed for each trial is set to its trial index.
+4.1.3 Evaluation Metrics
 In this study, we employ two evaluation metrics: the final simple regret and the area under the best‑so‑far regret curve (AUC).
@@
 Following this concept, the final simple regret is the regret value obtained in the last iteration and reflects the algorithm’s ultimate optimization capability when computational cost is disregarded.
-By contrast, the regret AUC—the sum (area under the curve) of the simple regret across all iterations—quantifies the convergence speed of the entire optimization process:
+By contrast, the regret AUC, or cumulative regret—the sum (area under the curve) of the simple regret across all iterations—quantifies the convergence speed of the entire optimization process:
 $\text{AUC}_{T}=\Sigma r_{t}$
-All experiments were executed for 200 iterations and repeated 20 times using an identical sequence of random seeds.
-Algorithm performance was evaluated by the mean and variance of the final simple regret and regret AUC on all 20 trials; in both cases, smaller values indicate better performance.
-Results and Discussion
-In this subsection, we present and discuss our experimental results along different dimensions against current SOTA algorithm BOPS-H.
-As shown in figure 1, the Merge kernel consistently outperforms the Mallows kernel across all four permutation optimization benchmarks.
-The regret curves demonstrate that MergeBO achieves faster convergence and lower mean regret compared to BOPS-H throughout the entire optimization process.
-Notably, the advantage is particularly pronounced on more challenging tasks such as QAP and Floor Planning, where the Merge kernel exhibits both reduced variance and a more rapid decrease in regret.
-The TSP problem is a relatively common and tractable combinatorial optimization problem, so both algorithms are able to approach the global optimum in the end.
-Nevertheless, it is evident that the Merge kernel achieves a faster convergence rate compared to the Mallows kernel.
-These results highlight the improved sample efficiency and robustness of the Merge kernel in complex permutation spaces.
-Table 1 provides a quantitative summary of performance, reporting both the final simple regret and the best-so-far area under the curve (AUC) for each benchmark.
-Merge kernel achieves either the lowest or statistically equivalent regret on all tasks, and outperforms Mallows kernel in three out of four benchmarks in terms of final regret and AUC.
-For instance, on QAP (n=15), the Merge kernel attains a final regret of $6578.0\pm 4579.2$, significantly better than the Mallows kernel’s $8090.4\pm 4101.9$, which can be attributed to the Merge kernel’s effective capture of both local and global ordering patterns inherent in facility-location scenarios.
-For Cell Placement task, Merge kernel’s slightly higher regret AUC but better stability indicates a robustness advantage, especially valuable in real-world circuit design scenarios.
-These results confirm that our proposed kernel offers superior optimization performance and generalizes effectively to a variety of real-world permutation problems.
-Ablation Study
-In this section, we perform ablation experiments to systematically evaluate how the exclusion of each individual trick affects the overall effectiveness of the kernel.
-The ablation study results in Table 2 demonstrate that each of the three auxiliary descriptors—pairwise comparison, pattern (sliding window motif), and shift histogram—contributes to the overall performance of the Merge kernel.
-When all tricks are included, the model consistently achieves either the best or nearly best results in both simple final regret and regret AUC across most benchmarks, highlighting the complementary effects of these modules.
-Removing any single component typically leads to a performance drop, underscoring the importance of each descriptor in capturing diverse structural features of permutation problems.
-We observe that removing the shift histogram module consistently results in a notable performance degradation across all tasks, as it provides a robust global positional signal, which is essential for quickly identifying globally optimal permutation structures, indicating that this component is the most impactful among the three descriptors in terms of optimization performance.
-Interestingly, the impact of the other two tricks varies by task.
-For instance, omitting the pairwise comparison module yields a lower final regret on the Cell Replacement task and the QAP task, however it also results to higher regret AUC, suggesting that local pairwise information may be a trade-off option for optimization speed and performance.
-In contrast, removing the pattern module leads to a noticeable decline in AUC for the Floor Planning benchmark, indicating its particular effectiveness in capturing local ordering relevant to that problem.
-Overall, these results confirm that the three lightweight descriptors are largely complementary, and that their joint inclusion ensures robust and superior optimization performance across a range of permutation-based tasks.
-Conclusions and Limitations
+Generally speaking, the final simple regret and AUC represent different, valuable aspects of an algorithm’s performance: the former represents the the quality of the solution it ultimately finds while the latter represents the “journey”, or convergence speed during optimization.
+For sample-efficient methods like BO, these two evaluation metrics are standard practices as they allow for a comprehensive comparison.
+Problem Dimension Merge feature length Mallows feature length
+TSP 15 45 105
+QAP 15 45 105
+FP 30 119 435
+CR 30 119 435
+TTP 280 2009 39060
+Table 1: Feature length comparison of Merge and Mallows kernel over problems of different scales.
+It is important to note that this study does not include a comparison of wall-clock computation times.
+This is a deliberate choice grounded in the fundamental premise of BO, where the cost of function evaluations (e.g., physical experiments or complex simulations) is assumed to far outweigh the computational cost of the algorithm itself.
+Consequently, our analysis prioritizes metrics related to sample efficiency, which is the primary bottleneck in such real-world scenarios.
+As a more stable and implementation-agnostic proxy for computational complexity, we instead report the feature vector dimensions generated by each kernel in Table 1, which directly reflects the compactness of the learned representations.
+Furthermore, as our experiments were conducted on a shared high-performance computing (HPC) cluster, reported wall-clock times would be subject to scheduler-induced variability, making them an unreliable metric for rigorous algorithmic comparison.
+4.2 Results and Discussion
+Figure 1: Low- and high-dimensional results comparing Mallows kernel (BOPS-H) and Merge kernel (MergeBO) on the current regret value (difference between best-so-far and optimal value) vs.
+number of iteration.
+Solid lines show the average regrets, while the shaded areas denote one standard deviation.
+Simple final regret Best so far AUC
+Merge Mallows Merge Mallows
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Simple final regret Best so far AUC
+Merge wins Ties Mallows wins Merge wins Ties Mallows wins
+$\text{TSP}_{n=15}$ 1 12 7 5 0 15
+$\text{QAP}_{n=15}$ 1 3 16 1 2 17
+$\text{FP}_{n=30}$ 10 4 6 10 1 9
+$\text{CR}_{n=30}$ 9 3 8 8 0 12
+$\text{TTP1}_{n=280}$ 50 0 0 50 0 0
+$\text{TTP2}_{n=280}$ 50 0 0 50 0 0
+$\text{TTP3}_{n=280}$ 50 0 0 50 0 0
+Table 2: Performance comparison between MergeBO and BOPS-H (Mallows kernel), bold font indicates statistically significant superiority as determined by a binomial sign test (p << 0.05, corresponding to more than 15 wins of 20 trials).
+Our experimental results, as presented in Table 2 and Figure 1, systematically reveal a strong correlation between the performance advantage of the Merge kernel and the problem’s dimensionality.
+For low-dimensional permutation problems (TSP and QAP), the Mallows kernel, which is specifically designed for such tasks, exhibited a slight performance advantage.
+However, on problems of intermediate dimensionality (FP and CR), the performance of the two kernels was comparable, with no statistically significant difference observed.
+This trend shifted decisively on the higher-dimensional TTP problems, where the Merge kernel demonstrated definitive superiority.
+The statistical data in Table 2 indicates that the Merge kernel significantly outperformed the Mallows kernel across all TTP instances on both the final regret and AUC metrics.
+Furthermore, the convergence curves in Figure 1 confirm its substantially faster convergence speed.
+This validates the superior performance and scalability of our proposed method on complex, high-dimensional optimization problems.
+These results corroborate our core hypothesis: the Merge kernel possesses an inherent advantage in high-dimensional permutation optimization problems, owing to its more compact structural design.
+The disparity between its $O(n\log n)$ feature complexity and the Mallows kernel’s $O(n^{2})$ complexity widens dramatically as the dimensionality n increases.
+This is explicitly quantified by the feature length comparison in Table 1: for low, intermediate, and high-dimensional problems, the feature dimensionality of the Mallows kernel is approximately 2, 3.5, and 19.5 times that of the Merge kernel, respectively.
+These results indicate that the experimental performance is governed by a trade-off between two key factors: (1) the ability to capture global information via the distance metric, and (2) the search efficiency driven by the compactness of the feature space.
+Due to its right-invariance property, the Mallows kernel possesses a stronger distance metric capability than the Merge kernel.
+However, as dimensionality increases, the vast disparity in feature vector length leads to a more pronounced space-compression effect.
+The resulting gains in search efficiency begin to outweigh the performance benefits afforded by the superior distance metric.
+Consequently, as the problem dimensionality continues to grow, the performance of the Merge kernel ultimately surpasses that of the Mallows kernel by a significant margin.
+This naturally suggests a dimension-dependent heuristic for practitioners: leveraging the Mallows kernel’s robust, right-invariant distance metric for low-dimensional tasks, while switching to the Merge kernel to capitalize on its superior scalability in high-dimensional regimes.
+5 Conclusions
 In this work, we proposed a novel kernel construction framework for permutation spaces by leveraging sorting algorithms as structured comparison schemes.
 Within this framework, we introduced the Merge kernel—an efficient, compact, and theoretically grounded alternative to the quadratic Mallows kernel.
-Through extensive experiments on both synthetic and real-world benchmarks, MergeBO consistently outperformed the state-of-the-art BOPS-H algorithm in terms of optimization performance, particularly on tasks where local structure and partial orderings dominate the objective landscape.
-Theoretically, we showed that the Merge kernel achieves the information-theoretic lower bound on feature complexity while preserving meaningful structure from the original permutation.
-This insight bridges sorting theory and kernel design, opening new directions for structured optimization.
-The proposed framework is modular and extensible, enabling the integration of other structured comparison strategies and additional descriptors without sacrificing computational efficiency.
-This opens possibilities for scaling BO to larger permutation spaces and more diverse application domains.
-Despite the promising results, our study still has several limitations.
-Our study isolates the impact of the Merge kernel by keeping BOPS-H’s neighbour-based local search and acquisition function unchanged.
-Consequently, most runtime is still dominated by neighbourhood exploration, so the $O(n\log n)$ feature map yields little speed-up on the modest problem sizes used here ($n\leq$30).
-Nor did we leverage gradient methods (e.g., L-BFGS-B) that could exploit the lower-dimensional feature space.
-In addition, our evaluation is confined to a single BO framework, leaving the kernel’s behaviour with alternative acquisition strategies or global search heuristics unexplored.
-Finally, we did not examine other fixed sorting–network algorithms such as bitonic sort because its $O(n\log n^{2})$ comparison count yields a feature vector that, for the problems with relatively small dimensions, can actually exceed the $O(n^{2})$ length of the Mallows kernel—thereby undermining our goal of dimensionality reduction.
-We therefore plan to (i) replace neighbour search with gradient-based acquisition optimization, (ii) applying the Merge kernel to AI tasks involving larger permutations, such as robotics planning or automated scientific discovery pipelines ($n>$100) to expose asymptotic gains, (iii) benchmark the Merge kernel within diverse BO frameworks to assess its generality, and (iv) generate and evaluate kernel function from bitonic sort algorithm on the larger permutation space.
-Acknowledgments
-We thank Xenophon Evangelopoulos from the University of Liverpool for drawing our attention to the lack of right‑invariance property in the Merge kernel, an insight that inspired us to devise lightweight descriptors to improve the algorithm’s performance.
+We showed that the Merge kernel achieves the information-theoretic lower bound on feature complexity while preserving meaningful structural information.
+This contribution bridges sorting theory and kernel design, revealing a fundamental trade-off between a model’s structural invariance and the compactness of its feature space.
+Empirical experiments confirmed this trade-off: while the state-of-the-art BOPS-H algorithm held a marginal advantage in low-dimensional problems, their performances were comparable on intermediate-dimensional tasks.
+In high-dimensional settings, however, MergeBO’s compact representation enabled far superior search efficiency, allowing it to significantly outperform BOPS-H.
+This work opens exciting possibilities for scaling Bayesian optimization to larger permutation spaces.
+Future directions include exploring other sorting algorithms with constant comparison counts, and applying this framework to challenging real-world domains like combinatorial neural architecture search and computational biology.
+Appendix A The Use of Large Language Models
+We declare that the large language models (LLMs) ChatGPT and Gemini are only used to aid and polish writing.
+No further applications of LLMs are used in this research, including but not limited to retrieval, research ideation and experiment designing.